In a right triangle, find the lenghth of the side not given.
b=2 c=sqrt 29 a=?

Answer by checkley77(12844)

a^2+2^2=sqrt29
a^2+4=5.385
a^2=5.385-4
a^2=1.385
a=sqrt1.385
a=1.177 answer.
proof:
1.177^2+4=5.385
1.385+4=5.385
5.385=5.385